# InterFormer: Towards Effective Heterogeneous Interaction Learning for Click-Through Rate Prediction

### v1

@@ 1. Introduction
Click-through rate (CTR) prediction, which aims to predict the probability of a user clicking an ad or item, is the fundamental task for various applications such as online advertising and recommender systems (Song et al., 2019; Lyu et al., 2020; Zhang et al., 2022, 2024). The quality of CTR prediction significantly influences the company revenue and user experience, drawing extensive attention from both academia and industry (Zhou et al., 2018, 2019; Lyu et al., 2020). For example, in online ad bidding, accurate CTR prediction helps advertisers optimize their bids and target the most receptive audiences. In content recommendation, accurate CTR prediction enables platforms to suggest more relevant content to users.
@@
Most of the existing CTR prediction models fall into two categories, including non-sequential models and sequential models. Non-sequential models focus on learning informative embeddings through feature interaction via inner-product (Lian et al., 2018; Sun et al., 2021), MLP (Wang et al., 2017, 2021b) and deep structured semantic model (Huang et al., 2013; Elkahky et al., 2015), but ignore the sequential information in user behaviors. Sequential models, in the contrast, employ additional modules, e.g., CNN (Tang and Wang, 2018), RNN (Sun et al., 2019; Zhou et al., 2018) and Attention modules (Lyu et al., 2020; Zhou et al., 2019; Zhai et al., 2024), to capture the sequential dependencies in user behaviors. Promising as it might be, existing sequential methods mostly employ a unidirectional information flow, where global information is used to guide sequence learning, while the reverse information flow from sequence to global information is largely ignored, hence suffering from insufficient inter-mode interaction. For example, global information often captures long-term interests, while sequence information reveals momentary interests, such as a sudden focus on a specific category of products, which can enhance the global context with immediate preference. Besides, due to the computational challenges of performing interaction learning among numerous global features and lengthy sequences, aggressive feature aggregation, e.g., sequence summation (Zhou et al., 2018), pooling (Xiao et al., 2020), and concatenation (Zhou et al., 2019), is often performed at early stages, inevitably leading to excessive information loss.
In light of the above limitations, we propose a novel heterogeneous interaction learning module named InterFormer, whose ideas are two-fold. To avoid insufficient inter-mode interaction, we enable bidirectional information flows between different modes, such that global and sequence learning are performed in an interleaving style. Specifically, to learn context-aware sequence embeddings, global sumamrization guides sequence learning via Personalized FFN (pFFN) and Multihead Attention (MHA) (Vaswani et al., 2017). To learn behavior-aware global embeddings, sequence summarization instructs global learning via an interaction module. To mitigate aggressive information aggregation, we adopt MHA for effective information selection, based on which the one-to-one mappings between input and output tokens are retained till the final interaction. Note that our framework is compatible with various interaction learning models like DCNv2 (Wang et al., 2021b), DHEN (Zhang et al., 2022), etc.
@@
  • Experiments and Analysis. We carry out extensive experiments on the proposed InterFormer with up to 0.14% AUC improvement on benchmark datasets and 0.15% Normalized Entropy (NE) gain on internal large-scale dataset. Besides, InterFormer exhibits promising scaling results in both feature scaling and model scaling.
@@ 3. Preliminaries
Table 1 summarizes the main symbols and notations used throughout the paper. We use bold uppercase letters for matrices (e.g., $\mathbf{X}$), bold lowercase letters for vectors (e.g., $\mathbf{x}$), and lowercase letters for scalars (e.g., $n$). The element at the $i$-th row and $j$-th column of a matrix $\mathbf{X}$ is denoted as $\mathbf{X}(i,j)$. The transpose of $\mathbf{X}$ is denoted by the superscript ${\scriptscriptstyle\mathsf{T}}$ (e.g., $\mathbf{X}^{\scriptscriptstyle\mathsf{T}}$). We use superscript $u$ to denote users, and subscripts $i$ and $t$ to denote item and timestamp, respectively (e.g., $y^{u}_{i_{t}}$). We use $\mathbf{x}_{i}^{(l)}$ to denote the $j$-th global feature of the $l$-th layer, and $\mathbf{s}_{t}^{(t)}$ to denote the sequence feature of the $l$-th layer at timestamp $t$. We consider the scenario with $m$ dense features, $n$ sparse features, and $k$ sequence features of length $T$. We use $d$ to denote the embedding dimension of the CTR model.
@@ 3.1. Click-Through Rate (CTR) Prediction
CTR prediction aims to estimate the probability of a user clicking on an item given heterogeneous information such as global context and behavior sequences. Formally, given a user set $\mathcal{U}$ and an item set $\mathcal{I}$, the interaction sequence of a user $u\in\mathcal{U}$ is defined as $S^{u}=[i^{u}_{1},i^{u}_{2},\dots,i^{u}_{T}]$, where $i^{u}_{t}\in\mathcal{I}$ is the item interacted at time step $t$. CTR prediction aims to estimate the probability of a user clicking on a new item $i^{u}_{T+1}$, denoted as $y^{u}_{i_{T+1}}$, given the historical interaction sequence $S^{u}$. Formally, we seek to learn a function $f:\mathcal{U}\times\mathcal{I}\times\mathcal{S}\rightarrow[0,1]$, where $\mathcal{S}$ is the set of all possible sequences, such that:
@@ 3.2. Feature Interaction
Inner Product-based Interaction. Given input feature vector $\mathbf{x}\in\mathbb{R}^{d}$, the inner product-based interaction, exemplified by Factorization Machines (FM) (Rendle, 2010), learns a latent vector $\mathbf{v}_{j}\in\mathbb{R}^{r}$ for each feature $j$, whose inner product $\langle\mathbf{v}_{j},\mathbf{v}_{k}\rangle$ describes the interaction strength between $\mathbf{x}(j)$ and $\mathbf{x}(k)$. Hence, the second-order feature interactions can be modeled as:
@@
where $\text{Norm}(\cdot)$ is a normalization function, $\text{Interaction}_{i}$ are different interaction modules that are further ensembled by $\text{Ensemble}_{i=1}^{k}(\cdot)$, e.g., summation and concatenation, and $\text{ShortCut}(\cdot)$ is an MLP serving as the residual connection for deep layer stacking.
@@ 4.1. Feature Preprocessing
In real-world scenarios, we should expect multiple sequences originating from different user actions (e.g., click and conversion), different platforms, etc. Besides, due to the uncertainty of user behaviors, these sequences often contain noisy and irrelevant user-item interactions. To this end, a MaskNet (Wang et al., 2021a) is employed to unify multiple sequences and filter out the internal noises via self-masking.
@@ 4.2. Global Arch: Towards Behavior-Aware Interaction Learning
where $\text{Interaction}^{(l)}(\cdot)$ is the interaction module. Note that we do not adhere to a specific interaction module, but rather, various backbone models, such as inner product, DCNv2 (Wang et al., 2021b) and DHEN (Zhang et al., 2022), can be adopted. Besides, an MLP is further employed to transform the output $\mathbf{X}^{(l+1)}$ to have consistent shape as the input $\mathbf{X}^{(l)}$ for selective information aggregation and easy layer-stacking.
By performing interaction learning on the concatenation of $\mathbf{X}^{(l)}$ and $\mathbf{S}^{(l)}_{\text{sum}}$, Eq. (7) incorporates three different interactions, including global-global, global-sequence, and sequence-sequence interactions. The global-global interactions capture explicit user interests by computing the relevance between the user profile and the target item content, while the global-sequence interactions capture implicit user interests by computing the relevance between the target item and the user’s current need behind the interaction sequence. Besides, by computing the interactions among sequence features, the most representative information can be promoted with high interaction scores, while inactive items with low scores can be filtered out. By stacking multiple layers, the Global Arch captures rich behavior-aware interactions at different orders.
@@ 4.3. Sequence Arch: Towards Context-Aware Sequence Modeling
In addition to the explicit static user preference in global features, the implicit dynamic user interests behind the behavior sequences provides complementary information (Zhou et al., 2019; Sun et al., 2019). However, given that user behavior sequence can be highly random and noisy, solely relying on sequence information is ineffective, and it is important to incorporate global context into sequence modeling. For example, a user may randomly browse items on online shopping platforms, but given the global information that the user is an electronic fan, electronic items, e.g., smartphone and laptops, in the browsing history are identified as key information require extra attention.
To learn context-aware sequence embeddings, a Sequence Arch is designed on top of two key ideas, including Personalized FFN (PFFN) and Multi-Head Attention (MHA). To enable interactions between global and sequence information, we first employ PFFN to transform sequence embeddings given global summarization as query. Given global summarization $\mathbf{X}_{\text{sum}}^{(l)}$ and sequence embedding $\mathbf{S}^{(l)}$ at the $l$-th layer, the PFFN operation is defined as:
@@
Besides, to model the relationship among events in a sequence, MHA is applied to enable the model to attend to different parts of the sequence, capturing long-range dependencies and contextual information. To incorporate global context, before feeding into the first InterFormer layer, the global summarization $\mathbf{X}^{(1)}_{\text{sum}}$ is prepended before the sequence as the CLS token, i.e., $\mathbf{S}^{(1)}=\left[\mathbf{X}^{(1)}_{\text{sum}}\|\mathbf{S}^{(1)}\right]$, hence the following MHA on CLS token can aggregate sequence information given global information as query (Vaswani et al., 2017; Devlin et al., 2018). This is similar to the early fusion idea described in Transact  (Xia et al., 2023) but we primarily consider ‘append’ instead of ‘concat’ as the fusion method. Besides, rotary position embeddings (Su et al., 2024) are applied on tokens so that the positional information in sequences can be effectively leveraged. In general, the Sequence Arch can be written as follow:
@@
Through layer stacking, for one thing, the sequence embeddings is aware of global interactions at different orders via PFFN; for another, MHA at different layers focuses on different parts of the sequence, capturing multi-scale sequence information. Therefore, the model can learn rich context-aware encoding of the sequential data capturing both local and global patterns within the sequence.
@@ 4.4. Bridging Arch: Towards Effective Information Selection and Summarization
For sequence information, three types of summarization are neatly designed, including CLS tokens, PMA tokens and recent interacted tokens. The CLS tokens $\mathbf{S}_{\text{CLS}}$ learned by MHA are selected as context-aware sequence summarization. However, the quality of CLS tokens largely depend on the learned global context. To compensate the heavy reliance on global context and enable more flexibility, the PMA tokens $\mathbf{S}_{\text{PMA}}$ (Lee et al., 2019), which are essentially sequence summarization based on learnable queries, are employed. Besides, the $K$ most recent interacted tokens $\mathbf{S}_{\text{recent}}$ are proven to be effective in capturing user’s recent interests (Borisyuk et al., 2024; Xia et al., 2023). The combination of the above information is further gated by a self-gating layer, serving as the behavior summarization for global interaction:
@@ 5.1. Experiment Setup
Baseline Methods. We compare the proposed InterFormer with 11 state-of-the-art models, including (1) non-sequential methods: FM (Rendle, 2010), xDeepFM (Lian et al., 2018), AutoInt+ (Song et al., 2019), DCNv2 (Wang et al., 2021b), FmFM (Sun et al., 2021), DOT product, DHEN (Zhang et al., 2022), Wukong (Zhang et al., 2024), and (2) sequential methods: DIN (Zhou et al., 2018), DIEN (Zhou et al., 2019), BST (Chen et al., 2019), DMIN (Xiao et al., 2020), DMR (Lyu et al., 2020). We adopt DHEN (Zhang et al., 2022) as the global arch for InterFormer in the experiments. Detailed model configurations are provided in Appendix B.2.
Metrics. We adopt three widely-used metrics to evaluate the models from different aspects, including:
@@
  • LogLoss (cross-entropy loss) measures the distance between the model prediction $\hat{y}$ and the actual label $y$, and can be computed as $L(y,\hat{y})=-\left(y\log\left(\hat{y}\right)+(1-y)\log\left(1-\hat{y}\right)\right)$. Lower the better.
@@ 5.2. Effectiveness Results
The experiment results are shown in Table 3. We first observe that sequential methods consistently outperforms non-sequential methods on all datasets. For non-sequential methods, sequence information is naively aggregated in early stages and further processed together with global information. While in sequential methods, neatly designed sequence processing modules, e.g., RNN (Zhou et al., 2019; Sun et al., 2019) and attention mechanism (Zhou et al., 2018; Chen et al., 2019; Xiao et al., 2017; Lyu et al., 2020), are employed, so that sequential information can be processed in aware of global context. The universal outperformance of sequential methods validates that different data modes should be processed differently. Besides, results show that aggressive sequence summarization in early stages in ignorance of global context will impair model performance.
Comparing InterFormer with other sequential methods, InterFormer achieves state-of-the-art performance. Specifically, InterFormer outperforms the best competitor by up to 0.14% in gAUC, 0.14% in AUC and 0.54% in LogLoss. These results demonstrate InterFormer’s effectiveness on diverse datasets and generalization across different recommendation tasks.

### v2

@@ 1. Introduction
Click-through rate (CTR) prediction, which aims to predict the probability of a user clicking an ad or item, is the fundamental task for various applications such as online advertising and recommender systems (Song et al., 2019; Lyu et al., 2020; Zhang et al., 2022, 2024). The quality of CTR prediction significantly influences the company revenue and user experience, drawing extensive attention from both academia and industry (Zhou et al., 2018, 2019). For example, in online ad bidding, accurate CTR prediction helps advertisers optimize their bids and target the most receptive audiences. In content recommendation, it enables platforms to suggest more relevant content to users.
@@
Most of the existing CTR prediction models fall into two categories, including non-sequential models and sequential models. Non-sequential models focus on learning informative embeddings through feature interaction via inner-product (Lian et al., 2018; Sun et al., 2021), MLP (Wang et al., 2017, 2021b) and deep structured semantic model (Huang et al., 2013; Elkahky et al., 2015), but ignore the sequential information in user behaviors. Sequential models, in contrast, employ additional modules like CNN (Tang and Wang, 2018), RNN (Sun et al., 2019; Zhou et al., 2018) and Attention modules (Lyu et al., 2020; Zhou et al., 2019; Zhai et al., 2024), to capture the sequential dependencies in user behaviors. Promising as it might be, existing sequential methods mostly employ a unidirectional information flow, where global information is used to guide sequence learning, while the reverse information flow from sequence to global information is largely ignored, resulting in insufficient inter-mode interaction. For example, global information often captures long-term interests, while sequence information reveals momentary interests, such as a sudden focus on a specific category of products, which can enhance the global context with immediate preference. Besides, due to the computational challenges of performing interaction learning among numerous global features and lengthy sequences, aggressive feature aggregation, e.g., sequence summation (Zhou et al., 2018), pooling (Xiao et al., 2020), and concatenation (Zhou et al., 2019), is often performed in the early stages, inevitably leading to excessive information loss.
In light of the above limitations, we propose a novel heterogeneous interaction learning module named InterFormer, whose ideas are two-fold. To avoid insufficient inter-mode interaction, we enable bidirectional information flows between different modes, such that global and sequence learning are performed in an interleaving style. Specifically, to learn context-aware sequence embeddings, global sumamrization guides sequence learning via Personalized FeedForward Network (PFFN) and Multihead Attention (MHA) (Vaswani et al., 2017). To learn behavior-aware global embeddings, sequence summarization instructs global learning via an interaction module. To mitigate aggressive information aggregation, we adopt MHA for effective information selection, based on which the one-to-one mappings between input and output tokens are retained till the final interaction. Note that our framework is compatible with various interaction learning models like DCNv2 (Wang et al., 2021b), DHEN (Zhang et al., 2022), etc.
@@
  • Experiments and Analysis. We carry out extensive experiments on the proposed InterFormer with up to 0.14% AUC improvement on benchmark datasets and 0.15% Normalized Entropy (NE) gain on internal large-scale dataset. In addition, we have conducted studies on internal deployment at Meta and find that InterFormer exhibits promising scaling results in both feature scaling and model scaling.
@@ 3. Preliminaries
$\odot$ Hadamard product
[…]
Table 1 summarizes the main symbols and notations used throughout the paper. We use bold uppercase letters for matrices (e.g., $\mathbf{X}$), bold lowercase letters for vectors (e.g., $\mathbf{x}$), and lowercase letters for scalars (e.g., $n$). The element at the $i$-th row and $j$-th column of a matrix $\mathbf{X}$ is denoted as $\mathbf{X}(i,j)$. The transpose of $\mathbf{X}$ is denoted by the superscript ${\scriptscriptstyle\mathsf{T}}$ (e.g., $\mathbf{X}^{\scriptscriptstyle\mathsf{T}}$). We use superscript $u$ to denote users, and subscripts $i$ and $t$ to denote item and timestamp, respectively (e.g., $y^{u}_{i_{t}}$). We use $\mathbf{x}_{j}^{(l)}$ to denote the $j$-th global feature of the $l$-th layer, and $\mathbf{s}_{t}^{(l)}$ to denote the sequence feature of the $l$-th layer at timestamp $t$. We consider the scenario with $m$ dense features, $n$ sparse features, and $k$ sequence features of length $T$. We use $d$ to denote the embedding dimension of the CTR model.
@@ 3.1. Click-Through Rate (CTR) Prediction
CTR prediction aims to estimate the probability of a user clicking on an item given heterogeneous information such as global context and behavior sequences. Formally, given a user set $\mathcal{U}$ and an item set $\mathcal{I}$, the interaction sequence of a user $u\in\mathcal{U}$ is defined as $S^{u}=[i^{u}_{1},i^{u}_{2},\dots,i^{u}_{T}]$, where $i^{u}_{t}\in\mathcal{I}$ is the item interacted at time step $t$ by user $u$. CTR prediction estimates the probability of a user clicking on a new item $i^{u}_{T+1}$, denoted as $y^{u}_{i_{T+1}}$, given the historical interaction sequence $S^{u}$. Formally, we seek to learn a function $f:\mathcal{U}\times\mathcal{I}\times\mathcal{S}\rightarrow[0,1]$, where $\mathcal{S}$ is the set of all possible sequences, such that:
@@ 3.2. Feature Interaction
Inner Product-based Interaction. Given input feature vector $\mathbf{x}\in\mathbb{R}^{d}$, the inner product-based interaction, exemplified by Factorization Machines (FM) (Rendle, 2010), learns a latent vector $\mathbf{v}_{j}\in\mathbb{R}^{r}$ for each feature $j$, whose inner product $\langle\mathbf{v}_{j},\mathbf{v}_{k}\rangle$ describes the interaction strength between features $\mathbf{x}(j)$ and $\mathbf{x}(k)$. Hence, the second-order feature interactions can be modeled as:
@@
where $\text{Norm}(\cdot)$ is a normalization function, $\text{Interaction}_{i}$ represents different interaction modules that are further assembled by $\text{Ensemble}_{i=1}^{k}(\cdot)$ such as summation and concatenation, and $\text{ShortCut}(\cdot)$ is an MLP serving as the residual connection for deep layer stacking.
@@ 4.1. Feature Preprocessing
Real-world scenarios often encounter multiple sequences originating from different user actions (e.g., click and conversion)or various platforms. Moreover, due to the uncertainty of user behaviors, these sequences often contain noisy and irrelevant user-item interactions. To address this, a MaskNet (Wang et al., 2021a) is employed to unify multiple sequences and filter out the internal noises via self-masking.
@@ 4.2. Global Arch: Towards Behavior-Aware Interaction Learning
where $\text{Interaction}^{(l)}(\cdot)$ is the interaction module. Note that we do not adhere to a specific interaction module, but rather, various backbone models, such as inner product, DCNv2 (Wang et al., 2021b) and DHEN (Zhang et al., 2022), can be adopted. Additionally, an MLP is used to transform the output $\mathbf{X}^{(l+1)}$ to match the shape of the input $\mathbf{X}^{(l)}$ enabling selective information aggregation and facilitating easy layer stacking.
By performing interaction learning on the concatenation of $\mathbf{X}^{(l)}$ and $\mathbf{S}^{(l)}_{\text{sum}}$, Eq. (7) incorporates three different interactions: global-global, global-sequence, and sequence-sequence interactions. The global-global interactions capture explicit user interests by computing the relevance between the user profile and the target item content, while the global-sequence interactions capture implicit user interests by computing the relevance between the target item and the user’s current need within the interaction sequence. Additionally, by computing the interactions among sequence features, the most representative information is promoted with high interaction scores, while inactive items with low scores are filtered out. By stacking multiple layers, the Global Arch captures rich behavior-aware interactions at different orders.
@@ 4.3. Sequence Arch: Towards Context-Aware Sequence Modeling
In addition to the explicit static user preference in global features, the implicit dynamic user interests behind the behavior sequences provides complementary information (Zhou et al., 2019; Sun et al., 2019). However, given that user behavior sequence can be highly random and noisy, solely relying on sequence information is ineffective. It is important to incorporate global context into sequence modeling. For example, a user may randomly browse items on online shopping platforms, but given the global information that the user is an electronic fan, electronic items, e.g., smartphone and laptops, in the browsing history are identified as key information that requires extra attention.
To learn context-aware sequence embeddings, a Sequence Arch is designed based on two key ideas: Personalized FFN (PFFN) and Multi-Head Attention (MHA). To enable interactions between global and sequence information, we first employ PFFN to transform sequence embeddings given global summarization as query. Given global summarization $\mathbf{X}_{\text{sum}}^{(l)}$ and sequence embedding $\mathbf{S}^{(l)}$ at the $l$-th layer, the PFFN operation is defined as:
@@
Besides, to model the relationship among events in a sequence, MHA is applied to enable the model to attend to different parts of the sequence, capturing long-range dependencies and contextual information. To incorporate global context, before feeding into the first InterFormer layer, the global summarization $\mathbf{X}^{(1)}_{\text{sum}}$ is prepended to the sequence as the CLS token, i.e., $\mathbf{S}^{(1)}=\left[\mathbf{X}^{(1)}_{\text{sum}}\|\mathbf{S}^{(1)}\right]$, hence the following MHA on the CLS token can aggregate sequence information using global information as the query (Vaswani et al., 2017; Devlin et al., 2018). This is similar to the early fusion idea described in Transact  (Xia et al., 2023) but we primarily consider ’append’ instead of ’concat’ as the fusion method. Additionally, rotary position embeddings (Su et al., 2024) are applied on tokens so that the positional information in sequences can be effectively leveraged. In general, the Sequence Arch can be written as follow:
@@
Through layer stacking, on one hand, the sequence embeddings is aware of global interactions at different orders via PFFN; on the other hand, MHA at different layers focuses on different parts of the sequence, capturing multi-scale sequence information. Therefore, the model can learn a rich context-aware encoding of the sequential data capturing both local and global patterns within the sequence.
@@ 4.4. Bridging Arch: Towards Effective Information Selection and Summarization
For sequence information, three types of summarization are neatly designed, including CLS tokens, PMA tokens and recent interacted tokens. The CLS tokens $\mathbf{S}_{\text{CLS}}$ learned by MHA are selected as context-aware sequence summarization. However, the quality of CLS tokens largely depend on the learned global context. To compensate for the heavy reliance on global context and enable more flexibility, the PMA tokens $\mathbf{S}_{\text{PMA}}$ (Lee et al., 2019), which are essentially sequence summarization based on learnable queries, are employed. Besides, the $K$ most recent interacted tokens ($\mathbf{S}_{\text{recent}}$) have been proven to be effective in capturing the user’s recent interests (Borisyuk et al., 2024; Xia et al., 2023). The combination of the above information is further gated by a self-gating layer, serving as the behavior summarization for global interaction:
@@ 5.1. Experiment Setup
Baseline Methods. We compare the proposed InterFormer with 11 state-of-the-art models, including (1) non-sequential methods: FM (Rendle, 2010), xDeepFM (Lian et al., 2018), AutoInt+ (Song et al., 2019), DCNv2 (Wang et al., 2021b), FmFM (Sun et al., 2021), DOT product, DHEN (Zhang et al., 2022), Wukong (Zhang et al., 2024), and (2) sequential methods: DIN (Zhou et al., 2018), DIEN (Zhou et al., 2019), BST (Chen et al., 2019), DMIN (Xiao et al., 2020), DMR (Lyu et al., 2020), TransAct (Xia et al., 2023). We adopt DHEN as the global arch for InterFormer in the experiments. Detailed model configurations are provided in Appendix B.2.
Metrics. We adopt four widely-used metrics to evaluate the models from different aspects, including:
@@
  • LogLoss (cross-entropy loss) measures the distance between the prediction $\hat{y}$ and the label $y$, and is computed as $L(y,\hat{y})\!=\!-\left(y\log\left(\hat{y}\right)+(1\!-\!y)\log\left(1\!-\!\hat{y}\right)\right)$. Lower the better.
@@ 5.2. Effectiveness Results
The experiment results are shown in Table 3. We first observe that sequential methods consistently outperforms non-sequential methods on all datasets. For non-sequential methods, sequence information is naively aggregated in early stages and further processed together with global information. While in sequential methods, neatly designed sequence processing modules, e.g., RNN (Zhou et al., 2019; Sun et al., 2019) and attention mechanism (Zhou et al., 2018; Chen et al., 2019; Xiao et al., 2017; Lyu et al., 2020), are employed, so that sequential information can be processed in aware of global context. The universal outperformance of sequential methods validates that different data modes should be processed differently. Besides, results show that aggressive sequence summarization in the early stages, without considering global context, will impair model performance.
Comparing InterFormer with other sequential methods, InterFormer achieves state-of-the-art performance. Specifically, InterFormer outperforms the best competitor by up to 0.9% in gAUC, 0.14% in AUC and 0.54% in LogLoss. These results demonstrate InterFormer’s effectiveness on diverse datasets and generalization across different recommendation tasks.